# Gravastar on the brane with a timelike extra dimension

### v1

@@ 2 Mathematical formalism and solutions
A modification of the Einstein–Hilbert action in the geometric sector of braneworld gravity leads to the emergence of additional contributions in the effective field equations on the brane. The most general form of the modified Einstein field equations (EFE) induced on the brane can be expressed as [52]
$m^{2}G_{\mu\nu}+\sigma h_{\mu\nu}=T_{\mu\nu}+\epsilon M^{3}\left(K_{\mu\nu}-h_{\mu\nu}K\right),$ (2.1)
@@
Employing the Gauss–Codazzi relations on the brane, the above equation can be recast into the form [52]
$G_{\mu\nu}+\Lambda_{\text{eff}}h_{\mu\nu}=8\pi G_{\text{eff}}T_{\mu\nu}+\frac{\epsilon}{1+\beta}\left[\frac{S_{\mu\nu}}{M^{6}}-W_{\mu\nu}\right].$ (2.2)
@@
$\beta=\frac{2\sigma m^{2}}{3M^{6}}.$ (2.3)
@@
For $\epsilon=+1$, the bulk spacetime has Lorentzian signature and the extra dimension is spacelike, corresponding to the Randall–Sundrum type II (RS II) model. In contrast, choosing $\epsilon=-1$ renders the extra dimension timelike, resulting in the Shtanov–Sahni (SS) braneworld scenario. The effective cosmological constant and gravitational coupling on the brane are given by
$\Lambda_{\text{eff}}=\frac{\Lambda_{\text{RS}}}{1+\beta},\qquad 8\pi G_{\text{eff}}=\frac{\beta}{m^{2}(1+\beta)},$ (2.5)
@@
$S_{\mu\nu}=\frac{1}{3}EE_{\mu\nu}-E_{\mu\alpha}E^{\alpha}_{\nu}+\frac{1}{2}\left(E_{\alpha\gamma}E^{\alpha\gamma}-\frac{1}{3}E^{2}\right)h_{\mu\nu},$ (2.7)
@@
The remaining correction term $W_{\mu\nu}$ arises from the projection of the bulk Weyl tensor $W_{\mu\nu\alpha\gamma}$ onto the brane and captures the influence of bulk graviton degrees of freedom. It is defined as
$W_{\mu\nu}=n^{\alpha}n^{\gamma}W_{\mu\nu\alpha\gamma}.$ (2.8)
@@
$8\pi G_{\text{eff}}=\frac{2\epsilon\sigma}{3M^{6}}.$ (2.10)
For the SS braneworld, where the extra dimension is timelike ($\epsilon=-1$), positivity of $G_{\text{eff}}$ requires the brane tension $\sigma$ to be negative. The effective cosmological constant on the brane in this limit is
$\Lambda_{\text{RS}}=\frac{\Lambda_{5}}{2}+\frac{\epsilon\sigma^{2}}{3M^{6}},$ (2.11)
@@
Following the original Randall-Sundrum construction, we impose the fine-tuning condition $\Lambda_{\text{RS}}=0$ in order to develop our wormhole model on the brane. Since the extra dimension is timelike, the evolution off the brane into the bulk constitutes a well-posed Cauchy problem with a valid solution in a neighborhood of the brane. Consequently, one may adopt an effective approach that focuses exclusively on the brane dynamics without specifying the global bulk geometry [54].
@@
$\frac{dp}{dr}=-\frac{1}{2}\frac{d\nu}{dr}(\rho+p).$ (2.12)
@@
$ds^{2}=-e^{\nu(r)}dt^{2}+e^{\lambda(r)}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (2.13)
From Eq. 2.1 we get
$e^{-\lambda}\left(\frac{\lambda^{\prime}}{r}-\frac{1}{r^{2}}\right)+\frac{1}{r^{2}}=\rho^{\mathrm{eff}},$ (2.14)
$e^{-\lambda}\left(\frac{\nu^{\prime}}{r}+\frac{1}{r^{2}}\right)-\frac{1}{r^{2}}=p_{r}^{\mathrm{eff}},$ (2.15)
$e^{-\lambda}\left(\frac{\nu^{\prime\prime}}{2}-\frac{\lambda^{\prime}\nu^{\prime}}{4}+\frac{\nu^{\prime 2}}{4}+\frac{\nu^{\prime}-\lambda^{\prime}}{2r}\right)=p_{t}^{\mathrm{eff}}.$ (2.16)
@@
The parameter $\rho=\rho_{c}$ is introduced as a constant critical energy density that characterizes the occurrence of a cosmological bounce. In the present framework, this critical density is related to the brane tension through
$\rho_{c}=2|\sigma|,$ (2.17)
where $\sigma$ denotes the brane tension. The absolute value is required since the Shtanov–Sahni (SS) braneworld scenario involves a negative brane tension.
The quantities $U$ and $P$ appearing in the modified Einstein field equations represent effective stress–energy components arising from the projection of the bulk Weyl tensor onto the brane. Specifically, $U$ corresponds to the effective energy density, while $P$ denotes the effective pressure induced on the brane due to bulk gravitational effects. A noteworthy feature of this construction is that, despite assuming ordinary matter on the brane to be a perfect fluid with isotropic pressure, the higher-dimensional contributions generate unequal effective radial and tangential pressures. This leads to an induced anisotropy in the effective pressure on the brane, quantified by a difference of magnitude $\frac{12P}{\rho_{c}}$, entirely originating from bulk effects. This observation lends support to the assertion made by Cattoën [53] that pressure anisotropy is a fundamental characteristic of gravastar configurations. While such anisotropy does not arise naturally within the standard relativistic description of perfect fluids, it emerges here as an inherent consequence of the underlying framework.
@@
$\tilde{P}_{\text{eff}}^{\text{avg}}=\frac{1}{3}\left(\tilde{P}^{r}_{\text{eff}}+2\tilde{P}^{t}_{\text{eff}}\right)=\frac{1}{3}\tilde{\rho}_{\text{eff}},$ (2.18)
@@
As discussed earlier, we adopt an effective brane-based approach [54] in which the detailed dynamics of the bulk spacetime are not explicitly specified. This allows considerable freedom in selecting the functional form of the projected Weyl quantities on the brane. Similar assumptions regarding the Weyl sector have been employed in [33, 55]. In the present analysis, we assume a linear equation of state relating the Weyl-induced pressure and energy density of the form
$P=\omega U,$ (2.21)
where $\omega$ is a constant parameter to be fixed using the junction conditions.
@@
For the spacetime geometry, the temporal metric component is assumed to follow the regular Kuchowicz ansatz [56], given by
$e^{\nu(r)}=e^{Br^{2}+2\ln C},$ (2.23)
where $B$ and $C$ are constant parameters. The constant $B$ carries dimensions of inverse length squared, whereas $C$ is a dimensionless quantity. The Kuchowicz potential is adopted primarily due to its regular and well-behaved nature over the entire range of finite radial coordinates. This property makes it particularly suitable for modeling the interior metric potential of non-singular compact configurations, such as gravastars and other regular collapse solutions [57, 58, 59]. Furthermore, the versatility of the Kuchowicz potential has been demonstrated in other gravitational settings, where it has successfully been employed as a redshift function in the construction of wormhole geometries [55].
2.1 Interior solution
@@
$p=-\rho=-\rho_{c},$ (2.25)
where $\rho_{c}$ denotes the constant energy density of the gravastar interior. Consequently, the pressure is also spatially uniform throughout the core.
@@
${{\rm e}^{-\lambda\left(r\right)}}={\frac{8{r}^{2}{{C_{1}}}^{2}\pi-8\rho_{\mathrm{c}}{C_{1}}\pi{r}^{2}-8{r}^{2}\left(\omega+1/2\right)\sqrt{\rho_{{0}}{C_{1}}}+\rho_{\mathrm{c}}}{2B{r}^{2}\rho_{\mathrm{c}}+\rho_{\mathrm{c}}}}$ (2.26)
As we see, it is found to depend on the critical energy density $\rho_{c}$ which puts an upper bound on the energy density contained in the brane, in terms of the fundamental physical parameter brane tension $\sigma$. In the high energy limit, the constant brane tension tends to acquire very small values, such that extra dimensional effects are significant. In the low energy limit $\sigma\rightarrow\infty$ and standard GR is recovered. Besides the brane tension, the metric potential also depends on the EoS parameter $\omega$, which basically describes an effective fictitious fluid on the brane, that basically models the non-local gravitational corrections arising from transfer of gravitational degrees of freedom from the bulk to the brane. There is also an additional dependence on the constant of integration $C_{1}$.
2.2 Active gravitational mass $M(R)$
@@
$M(R)$ $=$ $\int_{0}^{R_{1}}4\pi r^{2}\rho_{eff}dr,$ (2.27)
$=$ $\frac{4}{3}\pi R_{1}^{3}8\pi{C_{1}}\left(1-{\frac{{C_{1}}}{\rho_{\mathrm{c}}}}\right)-12{\frac{\sqrt{\rho_{{0}}{C_{1}}}}{\rho_{\mathrm{c}}}}$
@@
For the gravastar core, the equation of state $p=-\rho=-\rho_{c}$ implies a vacuum-like interior with constant energy density and negative pressure. In general relativity, the active gravitational mass density is governed by the combination $\rho+3p$, which for this equation of state becomes
$\rho+3p=-2\rho_{c}<0,$ (2.28)
@@
The explicit result reflects this repulsive character. The leading contribution, proportional to $R_{1}^{3}$, resembles the usual volume term of a constant-density configuration but is modified by the factor $\left(1-C_{1}/\rho_{c}\right)$, which suppresses the effective mass due to the negative pressure of the interior condensate. The second term enters with a negative sign and further reduces the total mass, arising from the high-energy brane corrections encapsulated in $\rho_{\mathrm{eff}}$.
Consequently, the total active gravitational mass of the gravastar interior can be strongly suppressed and may even become negative for suitable choices of the parameters $C_{1}$ and $\rho_{c}$. This negative or vanishing active mass provides a concrete dynamical realization of the repulsive nature of the gravastar core, preventing continued collapse and ensuring the absence of a central singularity. It is this feature that fundamentally distinguishes the gravastar interior from that of a classical black hole, even though both may share the same exterior geometry.
2.3 Intermediate Shell
@@
${{\rm e}^{-\lambda(r)}}={\frac{-24{r}^{2}{{C_{1}}}^{2}\pi{{\rm e}^{-2B{r}^{2}}}+8{{\rm e}^{-B{r}^{2}}}\pi{C_{1}}{r}^{2}\rho_{\mathrm{c}}-8{r}^{2}\left(\omega+1/2\right)\sqrt{\rho_{{0}}{C_{1}}}{{\rm e}^{-\frac{Br^{2}}{2}}}+\rho_{{c}}}{2B{r}^{2}\rho_{\mathrm{c}}+\rho_{\mathrm{c}}}},$ (2.29)
@@ 2.4 Exterior solution
$ds^{2}=-\left(1-\frac{2M}{r}-\frac{Q}{r^{2}}\right)dt^{2}+\left(1-\frac{2M}{r}-\frac{Q}{r^{2}}\right)^{-1}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$ (2.30)
@@
The influence of the bulk gravity is therefore not confined to the interior of the gravastar but extends into the exterior vacuum region as well. If the effective correction to the exterior energy density arising from the Weyl term is denoted by $U_{e}$, it scales inversely with the fourth power of the radial coordinate, such that $U_{e}\sim Q/r^{4}$. Consequently, the sign of the tidal charge is directly determined by the sign of $U_{e}$: a positive effective energy density corresponds to a positive tidal charge, while a negative $U_{e}$ leads to a negative value of $Q$.
In conventional braneworld models with a spacelike extra dimension, the effective energy density $U_{e}$ is typically negative in order to ensure the localization of gravity on the brane, which in turn implies a negative tidal charge. In contrast, the present analysis is carried out within a braneworld framework featuring a timelike extra dimension. Consistent with this choice, the positive branch of the effective energy density is adopted, both in the interior and in the exterior regions. The appearance of the negative sign in front of the tidal charge term in the metric is itself a direct consequence of the timelike nature of the extra dimension.
@@
3 Physical properties of the gravastar

### v2

@@ II Mathematical formalism and solutions
A modification of the Einstein–Hilbert action in the geometric sector of braneworld gravity leads to the emergence of additional contributions in the effective field equations on the brane. The most general form of the modified Einstein field equations (EFE) induced on the brane can be expressed as Sahni5
$m^{2}G_{\mu\nu}+\sigma h_{\mu\nu}=T_{\mu\nu}+\epsilon M^{3}\left(K_{\mu\nu}-h_{\mu\nu}K\right),$ (1)
@@
Employing the Gauss–Codazzi relations on the brane, the above equation can be recast into the form Sahni5
$G_{\mu\nu}+\Lambda_{\mathrm{eff}}h_{\mu\nu}=8\pi G_{\mathrm{eff}}T_{\mu\nu}+\frac{\epsilon}{1+\beta}\left[\frac{S_{\mu\nu}}{M^{6}}-W_{\mu\nu}\right].$ (2)
@@
$\beta=\frac{2\sigma m^{2}}{3M^{6}}.$ (3)
@@
For $\epsilon=+1$, the bulk spacetime has Lorentzian signature and the extra dimension is spacelike, corresponding to the Randall–Sundrum type II (RS II) model. In contrast, choosing $\epsilon=-1$ renders the extra dimension timelike, resulting in the SS (SS) braneworld scenario. The effective cosmological constant and gravitational coupling on the brane are given by
$\Lambda_{\mathrm{eff}}=\frac{\Lambda_{\text{RS}}}{1+\beta},\qquad 8\pi G_{\mathrm{eff}}=\frac{\beta}{m^{2}(1+\beta)},$ (5)
@@
$S_{\mu\nu}=\frac{1}{3}EE_{\mu\nu}-E_{\mu\alpha}e^{\alpha}_{\nu}+\frac{1}{2}\left(E_{\alpha\gamma}e^{\alpha\gamma}-\frac{1}{3}e^{2}\right)h_{\mu\nu},$ (7)
@@
The remaining correction term $W_{\mu\nu}$ arises from the projection of the bulk Weyl tensor $W_{\mu\nu\alpha\gamma}$ onto the brane and captures the influence of bulk graviton degrees of freedom. This quantity is defined as
$W_{\mu\nu}=n^{\alpha}n^{\gamma}W_{\mu\nu\alpha\gamma}.$ (8)
@@
$8\pi G_{\mathrm{eff}}=\frac{2\epsilon\sigma}{3M^{6}}.$ (10)
For the SS braneworld, where the extra dimension is timelike ($\epsilon=-1$), positivity of $G_{\mathrm{eff}}$ requires the brane tension $\sigma$ to be negative. The effective cosmological constant on the brane in this limit is
$\Lambda_{\text{RS}}=\frac{\Lambda_{5}}{2}+\frac{\epsilon\sigma^{2}}{3M^{6}},$ (11)
@@
Following the original Randall-Sundrum construction, we impose the fine-tuning condition $\Lambda_{\text{RS}}=0$ in order to develop our wormhole model on the brane. Since the extra dimension is timelike, the evolution off the brane into the bulk constitutes a well-posed Cauchy problem with a valid solution in a neighborhood of the brane. Consequently, one may adopt an effective approach that focuses exclusively on the brane dynamics without specifying the global bulk geometry S9 .
@@
$\frac{dp}{dr}=-\frac{1}{2}\frac{d\nu}{dr}(\rho+p).$ (12)
@@
$ds^{2}=-e^{\nu(r)}dt^{2}+e^{\lambda(r)}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (13)
It follows from Eq. (1) that
$e^{-\lambda}\left(\frac{\lambda^{\prime}}{r}-\frac{1}{r^{2}}\right)+\frac{1}{r^{2}}=\rho^{\mathrm{eff}},$ (14)
$e^{-\lambda}\left(\frac{\nu^{\prime}}{r}+\frac{1}{r^{2}}\right)-\frac{1}{r^{2}}=p_{r}^{\mathrm{eff}},$ (15)
$e^{-\lambda}\left(\frac{\nu^{\prime\prime}}{2}-\frac{\lambda^{\prime}\nu^{\prime}}{4}+\frac{\nu^{\prime 2}}{4}+\frac{\nu^{\prime}-\lambda^{\prime}}{2r}\right)=p_{t}^{\mathrm{eff}}.$ (16)
@@
The parameter $\rho=\rho_{\mathrm{c}}$ is introduced as a constant critical energy density that characterizes the occurrence of a cosmological bounce. In the present framework, this critical density is related to the brane tension through
$\rho_{\mathrm{c}}=2|\sigma|,$ (17)
where $\sigma$ denotes the brane tension. The absolute value is required since the SS (SS) braneworld scenario involves a negative brane tension.
The quantities $U$ and $P$ appearing in the modified Einstein field equations represent effective stress–energy components arising from the projection of the bulk Weyl tensor onto the brane. Specifically, $U$ corresponds to the effective energy density, while $P$ denotes the effective pressure induced on the brane due to bulk gravitational effects. A noteworthy feature of this construction is that, despite assuming ordinary matter on the brane to be a perfect fluid with isotropic pressure, the higher-dimensional contributions generate unequal effective radial and tangential pressures. This leads to an induced anisotropy in the effective pressure on the brane, quantified by a difference of magnitude $\frac{12P}{\rho_{\mathrm{c}}}$, entirely originating from bulk effects. This observation supports the viewpoint that pressure anisotropy is a fundamental characteristic of gravastar configurations Cattoen . While such anisotropy does not arise naturally within the standard relativistic description of perfect fluids, it emerges here as an inherent consequence of the underlying framework.
@@
$\tilde{P}_{\mathrm{eff}}^{\text{avg}}=\frac{1}{3}\left(\tilde{P}^{r}_{\mathrm{eff}}+2\tilde{P}^{t}_{\mathrm{eff}}\right)=\frac{1}{3}\tilde{\rho}_{\mathrm{eff}},$ (18)
@@
As discussed earlier, we adopt an effective brane-based approach S9 in which the detailed dynamics of the bulk spacetime are not explicitly specified. This allows considerable freedom in selecting the functional form of the projected Weyl quantities on the brane. Similar assumptions regarding the Weyl sector have been employed  Sengupta5 ; Sengupta2 . In the present analysis, we assume a linear equation of state relating the Weyl-induced pressure and energy density of the form
$P=wU,$ (21)
where $w$ is a constant parameter to be fixed using the junction conditions.
@@
For the spacetime geometry, the temporal metric component is assumed to follow the regular Kuchowicz ansatz Kuchowicz , given by
$e^{\nu(r)}=\exp({Br^{2}+2\ln C}),$ (23)
where $B$ and $C$ are constant parameters. The constant $B$ carries dimensions of inverse length squared, whereas $C$ is a dimensionless quantity. The Kuchowicz potential is adopted primarily due to its regular and well-behaved nature over the entire range of finite radial coordinates. This property makes it particularly suitable for modeling the interior metric potential of non-singular compact configurations, such as gravastars and other regular collapse solutions Biswas2020 . Furthermore, the versatility of the Kuchowicz potential has been demonstrated in other gravitational settings, where it has successfully been employed as a redshift function in the construction of wormhole geometries Sengupta2 .
II.1 Interior solution
@@
$p=-\rho=-\rho_{\mathrm{c}},$ (25)
where $\rho_{\mathrm{c}}$ denotes the constant energy density of the gravastar interior. Consequently, the pressure is also spatially uniform throughout the core.
@@
$e^{-\lambda(r)}=\frac{8\pi C_{1}^{2}r^{2}-8\pi\rho_{\mathrm{c}}C_{1}r^{2}-4r^{2}(2w+1)\sqrt{\rho_{0}C_{1}}+\rho_{\mathrm{c}}}{\rho_{\mathrm{c}}\left(2Br^{2}+1\right)}$ (26)
As we see, it is found to depend on the critical energy density $\rho_{\mathrm{c}}$ which puts an upper bound on the energy density contained in the brane, in terms of the fundamental physical parameter brane tension $\sigma$. In the high energy limit, the constant brane tension tends to acquire very small values, such that extra dimensional effects are significant. In the low energy limit $\sigma\rightarrow\infty$ and standard GR is recovered. Besides the brane tension, the metric potential also depends on the EoS parameter $w$, which basically describes an effective fictitious fluid on the brane, that basically models the non-local gravitational corrections arising from transfer of gravitational degrees of freedom from the bulk to the brane. There is also an additional dependence on the constant of integration $C_{1}$.
II.2 Active gravitational mass $M(R)$
@@
$M(R_{1})$ $=\int_{0}^{R_{1}}4\pi r^{2}\rho_{\mathrm{eff}}dr$
$=\frac{32}{3}\pi^{2}C_{1}R_{1}^{3}\left(1-\frac{C_{1}}{\rho_{\mathrm{c}}}\right)-\frac{12}{\rho_{\mathrm{c}}}\sqrt{\rho_{0}C_{1}}$ (27)
@@
For the gravastar core, the equation of state $p=-\rho=-\rho_{\mathrm{c}}$ implies a vacuum-like interior with constant energy density and negative pressure. In general relativity, the active gravitational mass density is governed by the combination $\rho+3p$, which for this equation of state becomes
$\rho+3p=-2\rho_{\mathrm{c}}<0,$ (28)
@@
The explicit result reflects this repulsive character. The leading contribution, proportional to $R_{1}^{3}$, resembles the usual volume term of a constant-density configuration but is modified by the factor $\left(1-C_{1}/\rho_{\mathrm{c}}\right)$, which suppresses the effective mass due to the negative pressure of the interior condensate. The second term enters with a negative sign and further reduces the total mass, arising from the high-energy brane corrections encapsulated in $\rho_{\mathrm{eff}}$.
Consequently, the total active gravitational mass of the gravastar interior can be strongly suppressed and may even become negative for suitable choices of the parameters $C_{1}$ and $\rho_{\mathrm{c}}$. This negative or vanishing active mass provides a concrete dynamical realization of the repulsive nature of the gravastar core, preventing continued collapse and ensuring the absence of a central singularity. It is this feature that fundamentally distinguishes the gravastar interior from that of a classical black hole, even though both may share the same exterior geometry.
II.3 Intermediate Shell
@@
$e^{-\lambda(r)}=\frac{-24\pi C_{1}^{2}r^{2}e^{-2Br^{2}}+{8\pi C_{1}r^{2}\rho_{\mathrm{c}}e^{-Br^{2}}}-4r^{2}(2w+1)\sqrt{\rho_{0}C_{1}}e^{-\frac{Br^{2}}{2}}+\rho_{\mathrm{c}}}{\rho_{\mathrm{c}}\left(2Br^{2}+1\right)}.$ (29)
@@ II.4 Exterior solution
$ds^{2}=-\left(1-\frac{2M}{r}-\frac{Q}{r^{2}}\right)dt^{2}+\left(1-\frac{2M}{r}-\frac{Q}{r^{2}}\right)^{-1}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$ (30)
@@
The influence of the bulk gravity is therefore not confined to the interior of the gravastar but extends into the exterior vacuum region as well. If the effective correction to the exterior energy density arising from the Weyl term is denoted by $U_{\text{e}}$, it scales inversely with the fourth power of the radial coordinate, such that $U_{\text{e}}\sim Q/r^{4}$. Consequently, the sign of the tidal charge is directly determined by the sign of $U_{\text{e}}$: a positive effective energy density corresponds to a positive tidal charge, while a negative $U_{\text{e}}$ leads to a negative value of $Q$.
In conventional braneworld models with a spacelike extra dimension, the effective energy density $U_{\text{e}}$ is typically negative in order to ensure the localization of gravity on the brane, which in turn implies a negative tidal charge. In contrast, the present analysis is carried out within a braneworld framework featuring a timelike extra dimension. Consistent with this choice, the positive branch of the effective energy density is adopted, both in the interior and in the exterior regions. The appearance of the negative sign in front of the tidal charge term in the metric is itself a direct consequence of the timelike nature of the extra dimension.
@@
III Physical properties of the gravastar